OpenStudy (yisselluna):

Which of the following would not be the result of a substitution in the following system? 2x + y = 7 y - x = 1 7 - 2x - x = 1 2x + x + 1 = 7 2(y + 1) + y = 7

OpenStudy (jiteshmeghwal9):

2x+y=7 y-x=1 from 2nd equation x=y-1...(3) put this third equation in equation (1)

OpenStudy (jiteshmeghwal9):

y=1+x...(4) put this 4th equation in 2nd equation

OpenStudy (jiteshmeghwal9):

finally check